Find the differential equation of the space curve in which the two families of
surfaces
u = x^2 − y^2 = c1 and
v = y^2 − z^2 = c2 intersect.
1
Expert's answer
2020-03-27T12:53:09-0400

"u = x^2 − y^2 = c_1...(1) \newline

v = y^2 − z^2 = c2 ...(2)"

differentiating both (1) and (2)

"du=2xdx-2ydy=0\newline
\implies xdx=ydy ...(3)\newline
dv=2ydy-2zdz=0\newline
\implies ydy=zdz...(4)
\newline
by\hspace{0.1cm} (3) \hspace{0.1cm}and \hspace{0.1cm} (4)\newline\implies xdx=ydy=zdz 
\newline
\implies \frac{xdx}{xyz}=\frac{ydy}{xyz}=\frac{zdz}{xyz}\newline
\implies \frac{dx}{yz}= \frac{dy}{xz}= \frac{dz}{xy}"

Answer:"\frac{dx}{yz}= \frac{dy}{xz}= \frac{dz}{xy}"